Monochromatic light with wavelength of 750 nm passes through two narrow slits. If the slits are( 25 x10-6 m) apart, at what angle will the fourth order bright fringe appear on a viewing screen?
1
Expert's answer
2020-05-15T16:20:03-0400

The angle at which the m-th order bright fringe occurs


"\sin\theta=\frac{m\lambda}{d}"

"\sin\theta=\frac{4\times 750\times 10^{-9}\:\rm m}{25\times 10^{-6}\:\rm m}=0.12"

"\theta=6.89^{\circ}"